OpenStudy (anonymous):

what is the spuare rout of 1/100

myininaya (myininaya):

\[\sqrt{1}=1 and \sqrt{100}=10\]

myininaya (myininaya):

\[\sqrt{\frac{1}{100}}=\frac{\sqrt{1}}{\sqrt{100}}\]

OpenStudy (anonymous):

1/10